mistyyrosee03:

If the measure of AC = 75, what are the lengths of AB and BC?

mistyyrosee03:

1 attachment
Karla:

AC=5x+11x-5. That is 16x-5 As AC=75 So 16x-5=75 That is 16x=75+5 16x=80 X=80/16 x=5 Put the value of x for AB and BC As AB=5x AB=5*5=25 BC=11(5)-5 BC=55-5 BC=50

mistyyrosee03:

Thank you